Calculate the molar mass of H2S04 ? (the unit for molar mass is g/mol) 



1
Expert's answer
2020-10-08T13:54:30-0400

"M (H2SO4)=2 A(H) +A(S)+4 A(O)= 2 +32+4 *16=98"